OpenStudy (anonymous):

idk if my brain is just turning off because it's late... but would someone please help me with solving these two equations? \[4^x + 6(4^{-x}) =5\] \[8(5^{2x}) + 8 (5^x) = 6\]

OpenStudy (anonymous):

if you multiply the first one by 4^x you get a quadratic in 4^x

OpenStudy (anonymous):

Hint: for the first equation, a quadratic in 4^x

OpenStudy (anonymous):

second one is a similar idea, but it's already 'cleaned up' for you

OpenStudy (anonymous):

e u=4^x solve u^2 -5u +6 =0 for u... then use that in u=4^x to find x (fixed a sign)

OpenStudy (anonymous):

Let u be 4^x \[u + 6u^{-1} = 5\]Multiply both sides by u.\[u(u + 6u^{-1}) = 5u\]\[u^2 + 6 = 5u\]\[u^2 - 5u + 6 = 0\] Use quadratic formula to solve for u.

OpenStudy (anonymous):

@alexeis_nicole ?

OpenStudy (anonymous):

hmm. sooo. when i solved using the quadratic formula i got u = 3, 2 @micahwood50

OpenStudy (anonymous):

it's factorable actually, but yeah.

OpenStudy (anonymous):

Correct. Since we now know that\[u = 4^x\] Now plug in values then find the x.

OpenStudy (anonymous):

so x = 64, or 16 ?

OpenStudy (anonymous):

3=4^x , 2=4^x ln3/ln4 =x, ln2/ln4 =x

OpenStudy (anonymous):

oooh wait

OpenStudy (anonymous):

whooops my bad

OpenStudy (anonymous):

OR: \[\Large x = \log_4 3 \text{ and }\log_4 2 \text{ or } \frac{1}{2}\]

OpenStudy (anonymous):

its on 0.79 and 0.5

OpenStudy (anonymous):

Yeah. That's right.

OpenStudy (anonymous):

Is this clear? Can you solve second problem?